A ray of light hits a glass pane with 60o angle of incidence. The pane is 5 mm thick, with index of refraction 1.54. Light is reflected by a mirror that is in contact with the back of the pane. By how much is the beam laterally displaced with respect to the return path if the pane is removed?
1
Expert's answer
2020-04-30T10:34:22-0400

As per the question,

Angle of incidence of the light beam"(i)=60^\circ"

Thickness of the pane "(t)=5mm =5\times 10^{-3}m"

refractive index of the pane "(n)=1.54"

Let the angle of refraction is =r,

"\sin (i)/\sin(r)=\mu"

"\Rightarrow \sin (r)=\dfrac{\sin(60^\circ)}{1.54}=\dfrac{\sqrt(3)}{2\times 1.54}=34^\circ"

Now, lateral displacement of the beam of light "=\dfrac{t}{\cos(r)}\sin(i-r)"

"=\dfrac{5\times 10^{-3}\sin(60^\circ-34^\circ)}{\cos(34^\circ)}=2.6mm"